Consider the value of t such that the area under the curve between −|t| and |t| equals 0.9.

Step 2 of 2 : Assuming the degrees of freedom equals 11, select the t value from the t table.
Answer #1

Solution:-

Given that,

degrees of freedom = 11

Using t table,

P( -t < T < t) = 0.9

= P(T < t) - P(T <-t ) = 0.9

= 2P(T < t) - 1 = 0.9

= 2P(T < t) = 1 + 0.9

= P(T < t) = 1.9 / 2

= P(T < t) = 0.95

= P(T < 1.796) = 0.95

= t ± 1.796